# Characterisation of Open Quantum System Dynamics based on Information Back-flow

### v1

@@ Abstract
For unital dynamics, we show that a generalized trace distance measure offers no advantage over the trace distance measure for witnessing non-Markovianity. We determine the class of non-unital channels where the standard trace distance measure is insufficient here and the generalized measure is necessary. Finally, we assess the status of the GTD measure as an indicator of information flow between an open system and its environment.
@@ I Introduction
The isolation of a quantum system from its environment is an ideal condition, rendering open systems a significant object of study Breuer and Petruccione (2007). Such investigations are involved in achieving quantum control, focusing on the analysis of complete positive channels (CP) Choi (1975); Sudarshan et al. (1961). CP maps distinguish themselves from positive (P) maps Choi (1972) by remaining unaffected by including ancillary reference systems that do not participate during evolution. Any uncorrelated system and the environment with combined Hamiltonian evolution guarantees the CP evolution. Although certain quantum-correlated Shabani and Lidar (2009); Jagadish et al. (2023) states may induce CP dynamics, such occurrences are not universal across all unitary dynamics. The correlations generally lead to various dynamical maps, not just CP maps Jordan et al. (2004); Pechukas (1994); Štelmachovič and Bužek (2001); Colla et al. (2022); Dijkstra and Tanimura (2012).
From start to finish, the channel goes through different temporal dynamics depending on various possible environments and their interaction with the system. This time evolution determines different processes based on these interactions. Should the environment be sufficiently big or the interaction remains sufficiently weak, it is reasonable to assume that the environment does not vary too much alongside the system and that the correlations are not significant enough to be considered. Under such circumstances, the time evolution is laid out by the Gorini-Kossakowski-Sudarshan-Lindblad (GKSL) Gorini et al. (1976); Lindblad (1976) time-local equation, which follows the semi-group property characteristic of Markovian evolution. As quantum technology advances, allowing for experimentally greater control of the system’s dynamics, the above assumptions can be relaxed White et al. (2020); Bernardes et al. (2015); Gulati et al. (2024); Liu et al. (2013). In that situation, semi-group dynamics will be replaced by general divisible dynamics, provided the dynamics is invertible.
@@
Breuer-Laine-Piilo Laine et al. (2010); Breuer et al. (2009) introduced an information-theoretic measure called BLP to understand the memory effect Breuer et al. (2016); Liu et al. (2013). It is based on the distinguishability of two quantum states, as the trace distance measure indicates. The CP and P processes weaken distinguishability, suggesting a loss of information to the environment. Since BLP does not respond to the translation part of the dynamics, a generalized trace distance measure (GBLP) was introduced. Here, the increment of generalized trace distance for finite time is identified with non-Markovianity. In an attempt to unify the non-Markovianity based on trace distance and divisibility, a theorem in Ref. Chruściński et al. (2018) shows that P-divisibility is connected to a trace distance condition, specifically the GBLP condition Chruściński et al. (2011); Wißmann et al. (2015) rather than the BLP condition. This would raise the hope that the GBLP measure should serve as the right indicator of the information flow Chruściński et al. (2018, 2011); Wißmann et al. (2015). However, here, we argue that it may not be so.
The paper is written in the following way: In II, the affine transformation of a qubit is defined, and the BLP and GBLP measures are introduced to characterize non-Markovianity. An example is shown for which BLP fails, but GBLP works. In Sec. III, the general class of dynamics where GBLP is necessary is determined. In Sec. IV, it is shown that GBLP is not the right indicator, in general, to capture the information flow. Finally, we conclude with discussions and conclusions in Sec. V.
@@ II.1 BLP measure
Breuer-Laine-Piilo Laine et al. (2010); Breuer et al. (2009) defined a measure that connects non-Markovianity with the information back-flow. This measure is important because, through this, memory dependence on dynamics was highlighted, which was not possible through divisibility-based non-Markovianity. It is not necessary to know the dynamical map to witness the non-Markovian behaviour of the dynamics experimentally. For defining the measure, they considered the trace distance defined as
@@
here, $P_{\rm dis}$ is the probability to distinguish two states. Unitary dynamics keeps this distance unchanged for all initial pairs; hence, distinguishability is unchanged under unitary dynamics. This is consistent with the fact that the system does not interact with any other quantum system under unitary dynamics, so there is no exchange of information. In the event of interactions, based on distinguishability, there are two categories of dynamics: (a) Markovian dynamics: the distance does not increase for any initial pair for any interval of time. In this case, there is never an increase of distinguishability. (b) Non-Markovian dynamics: the distance increases at least for a few initial pairs for some time interval. In this case, the distinguishability may increase for an interval of time.
CP-divisible dynamics is CP for all intermediate-initial to intermediate-final times; this will keep contracting for all intermediate times since CP maps contract. P-divisible dynamics is CP from initial to final time, but for some intermediate time, it can be just P, not CP; it will also keep contracting because P-maps also contract. Because of this, CP-divisible and P-divisible dynamics are considered Markovian according to this measure Chruściński et al. (2018).
@@ II.2 Generalized measure
Thus, the BLP definition does not respond to the translation part of the dynamics. If two maps have the same $T$ part with different $\vec{c}$, then BLP can not differentiate them. To overcome this, the GBLP measure, discussed below, was introduced. The Helstorm Matrix $\Delta_{p}(\rho_{1},\rho_{2})=p\rho_{1}-q\rho_{2}$ with $p+q=1$ was considered for the distance to overcome this issue because it is sensitive to translations Chruściński et al. (2011). The generalized distance is defined as
@@
which is also a distance measure between two states but prepared with biased probabilities. This is also shown by connecting the distinguishability of two states with the generalized distance Chruściński et al. (2011); Wißmann et al. (2015). Unlike the BLP measure, P-divisibility is known to be connected with this generalised measure with the if and only if condition.
As a simple example where BLP fails but GBLP works, consider the non-P-divisible generalized amplitude damping noise given by the following affine map $\mathcal{E}(\vec{r})=T(t)\vec{r}+\vec{c}(t)$, where
$T(t)$ $=\begin{pmatrix}\sqrt{\eta}&0&0\\ 0&\sqrt{\eta}&0\\ 0&0&\eta\end{pmatrix}$
$\vec{c}(t)$ $=\begin{pmatrix}0\\ 0\\ (2s-1)(1-\eta)\end{pmatrix}$ (5)
where, $\eta=e^{-\gamma t}$ for $\gamma>0$ and $s=\cos^{2}{ft}$ for all $f$. Consider the initial states $\vec{r}_{1}\equiv(x_{1},y_{1},z_{1})^{\dagger}$ and $\vec{r}_{2}\equiv(x_{2},y_{2},z_{2})^{\dagger}$. We have the trace distance between their evolved versions to be
$\frac{1}{2}\textrm{Tr}[\mathcal{E}_{t}(\rho_{1})-\mathcal{E}_{t}(\rho_{2})]$ $=\frac{1}{2}|T(\textbf{r}_{1})-T(\textbf{r}_{2})|$
@@
$\textrm{Tr}[p\mathcal{E}_{t}(\rho_{1})-q\mathcal{E}_{t}(\rho_{2})]$ $\neq|pT(\vec{r}_{1})-qT(\vec{r}_{2})+(p-q)\vec{c}|$ (7)
@@
Figure 1: Generalised distance between states $\dfrac{1}{2}(\mathbb{I}+\sigma_{z})$ and $\dfrac{1}{2}(\mathbb{I}+\sigma_{x})$ for $p=0.25$ is plotted as a function of time for generalised amplitude damping noise (See Eq. II.2). Noise parameters $\gamma$ and $f$ are 0.1 and 4, respectively. Distance and time are in arbitrary units.
@@ III Unitality, divisibility and distance
The above represents one example where the two measures differ, but we want to determine the general classes where they differ. We are checking the distance evolution for both BLP and GBLP measures for CP, P and non-P maps. Once that is understood, the connection with the whole dynamics is straightforward because, for divisible dynamics, the whole dynamics is nothing but the composition of the intermediate maps. The GBLP measure is calculated for two initial states as $\rho_{i}\equiv(x_{i},y_{i},z_{i})$, where $i=1,2$ and probability $p$. The eigenvalues of $\Delta^{\dagger}\Delta$ are
@@
To calculate the distance, we add the square root of the eigenvalues because the trace does not depend on the basis. The generalised distance is
@@
$D(t)=\begin{array}[]{ll}\lvert\vec{w}(t)\rvert,&\lvert\vec{w}(t)\rvert\geq\lvert p-q\rvert\\ p-q,&\lvert\vec{w}(t)\rvert<\lvert p-q\rvert\\ \end{array}$ (8)
We notice that if $p=q$, the GBLP distance reverts to the BLP distance.
@@
The GBLP and BLP conditions are equivalent for unital dynamics.
@@
The translation vector vanishes for unital dynamics, and the dynamics is characterized by rotation, compression or expansion. Eq. (8) reduces to:
$D(t)=\begin{array}[]{ll}\lvert T(p\vec{r}_{1}-q\vec{r}_{2})\rvert,&\lvert\vec{w}(t)\rvert\geq\lvert p-q\rvert\\ p-q,&\lvert\vec{w}(t)\rvert<\lvert p-q\rvert\\ \end{array}$ (9)
@@
    If the channel is GBLP NM, then $\lvert T(p\vec{r}_{1}-q\vec{r}_{2})\rvert>\lvert p\vec{r}_{1}-q\vec{r}_{2}\rvert$ for some $p$, $\vec{r}_{1}$, $\vec{r}_{2}$, and $t$. Since $p,q\leq 1$, $p\vec{r}_{1}$ and $q\vec{r}_{2}$ will also be valid vectors belonging to the Bloch sphere. Hence $\lvert T(\vec{r^{\prime}}_{1}-\vec{r^{\prime}}_{2})\rvert>\lvert\vec{r^{\prime}}_{1}-\vec{r^{\prime}}_{2}\rvert$ where $\vec{r^{\prime}}_{1}=p\vec{r}_{1}$ and $\vec{r^{\prime}}_{2}=q\vec{r}_{2}$. This is the condition for BLP-NM.
@@
    If BLP is NM, $\lvert T(\vec{r}_{1}-\vec{r}_{2})\rvert>\lvert\vec{r}_{1}-\vec{r}_{2}\rvert$ for some $\vec{r}_{1}$ and $\vec{r}_{2}$. This condition gives one pair that belongs to GBLP with $p=\dfrac{1}{2}=q$. That makes GBLP also NM.
As a corollary of Theorem 1, there is no unital non-P-divisible BLP Markovian dynamics. We now consider the question of what are the channels for which the BLP and GBLP criteria are inequivalent.
@@
BLP and GBLP differ only for non-unital non-P-divisible BLP dynamics.
@@
Our earlier example showed that the BLP and GBLP conditions are inequivalent. From Theorem 1, it follows that, nevertheless, they are equivalent for unital noise. Thus, their inequivalence can hold only for non-unital noise. Here, by Ref. Chruściński et al. (2018), we know that P-divisibility is equivalent to the GBLP Markovianity. Thus, the inequivalence of the BLP and GBLP conditions can arise only from non-unital non-P-divisible noise. If this noise is BLP non-Markovian, then the two criteria will agree. ∎

### v2

# On the distinction between distinguishability of states and witness of non-Markovianity

@@ Abstract
Non-P-divisibility is the strongest divisibility-based notion of quantum non-Markovianity. The generalized trace distance (GTD) based criterion is known to be an optimal witness of non-P-divisibility of dynamical maps, in the sense that a given map is non-P-divisible if and only if there exists a pair of states that demonstrates increased distinguishability in the GTD sense. This observation forms the basis for associating an information backflow with this type of non-Markovianity. Nevertheless, we point out that the criterion is not a tight witness of distinguishability when applied to individual pairs of states; specifically, there can exist a pair of states whose distinguishability manifests in an increase, but the GTD criterion fails to indicate this. Note that this is beside the fact that pairs of states can exist whose distinguishability doesn’t evolve under a non-P-divisible map. Worse, the GTD criterion is not a faithful witness in that manifestly indistinguishable states that are indicated to be GTD distinguishable can exist. In other words, the case of GTD shows that a witness of non-Markovianity is not necessarily a universally applicable witness of distinguishability or information backflow across the entire state space. Furthermore, we demonstrate that for qubit unital dynamics, the GTD-based measure provides no advantage over the standard trace distance measure in witnessing non-Markovianity. We determine the class of qubit non-unital channels where the standard trace distance measure is insufficient and the generalized measure is necessary.
@@ I Introduction
The isolation of a quantum system from its environment is an ideal condition, rendering open systems a significant object of study Breuer and Petruccione (2007). Such investigations are involved in achieving quantum control, focusing on the analysis of complete positive channels (CP) Choi (1975); Sudarshan et al. (1961). CP maps distinguish themselves from positive (P) maps Choi (1972) by remaining unaffected by including ancillary reference systems that do not participate during evolution. Any uncorrelated system and the environment with a combined Hamiltonian evolution guarantees the CP evolution. Although certain quantum-correlated Shabani and Lidar (2009); Jagadish et al. (2023) states may induce CP dynamics, such occurrences are not universal across all unitary dynamics. The correlations generally lead to various dynamical maps, not just CP maps Jordan et al. (2004); Pechukas (1994); Štelmachovič and Bužek (2001); Colla et al. (2022); Dijkstra and Tanimura (2012).
From start to finish, the channel goes through different temporal dynamics depending on various possible environments and their interaction with the system. This time evolution determines different processes based on these interactions. Should the environment be sufficiently large or the interaction remain sufficiently weak, it is reasonable to assume that the environment does not vary significantly alongside the system and that the correlations are not strong enough to be considered significant. Under such circumstances, the time evolution is laid out by the Gorini-Kossakowski-Sudarshan-Lindblad (GKSL) Gorini et al. (1976); Lindblad (1976) time-local equation, which follows the semi-group property characteristic of Markovian evolution. As quantum technology advances, allowing for experimentally greater control of the system’s dynamics, the above assumptions can be relaxed White et al. (2020); Bernardes et al. (2015); Gulati et al. (2024); Liu et al. (2013a). In that situation, semi-group dynamics will be replaced by general divisible dynamics, provided the dynamics is invertible.
@@
Breuer-Laine-Piilo Laine et al. (2010); Breuer et al. (2009) introduced an information-theoretic measure called BLP to understand the memory effect Breuer et al. (2016); Liu et al. (2013a). It is based on the distinguishability of two quantum states, as indicated by the trace distance measure. The CP and P processes weaken distinguishability, suggesting a loss of information to the environment. Since BLP does not respond to the translation part of the dynamics Liu et al. (2013b), a generalized trace distance measure (GBLP) was introduced Chruściński et al. (2011). Here, the increment of the generalized trace distance for finite time is identified with non-Markovianity. In an attempt to unify the non-Markovianity based on trace distance and divisibility, a theorem in Ref. Chruściński et al. (2018) shows that P-divisibility is connected to a trace distance condition, specifically the GBLP condition Chruściński et al. (2011); Wißmann et al. (2015) rather than the BLP condition, i.e., an invertible dynamical map is P-divisible if and only if a monotonic non-increase of GTD distinguishability is observed for any pair of states of the system in question. The GBLP condition thus serves as an optimal witness of the non-P-divisibility of a map. In related work, Ref. Bylicka et al. (2017) shows that for an invertible dynamical map, CP-divisibility is equivalent to a monotonic non-increase of distinguishability for two equiprobable states of the system together with an ancilla, giving an explicit method to construct a pair of such states in the case of a non-CP-divisible map.
[…]
Even so, we expect that the concept of distinguishability or information flow must be applicable even at the microscopic level, i.e., at the level of pairs of individual evolving states. Any pair of states whose distinguishability manifestly evolves must be witnessed as such by the GTD criterion and vice versa. In this context, a tight witness is one that doesn’t admit false negatives, i.e., all positive cases of outflow or inflow are identified as such and not flagged as no-flow. On the other hand, a faithful witness doesn’t admit false positives (only positive cases are indicated as such). We may accept a non-tight witness, but an unfaithful witness would merit caution.
This raises the question of whether the GTD criterion can serve as a faithful witness to distinguishability or information flow when considering individual pairs of states. Here, we will find that this is not the case: it is non-tight in that there can be a pair of states with evolving distinguishability, but the GTD criterion does not indicate this. More worryingly, it is not faithful in that a pair of states can be manifestly undistinguishable but shown to be GTD distinguishable. In other words, we will find that the GTD criterion, based on which GBLP is neither a tight nor even a faithful witness of information backflow.
The paper is written in the following way: In II, the affine transformation of a qubit is defined, and the BLP and GBLP measures are introduced to characterise non-Markovianity. An example is shown for which BLP fails, but GBLP works. In Sec. III, the general class of dynamics where GBLP is necessary is determined. In Sec. IV, it is shown that GBLP is not the right indicator, in general, to capture the information flow. Finally, we conclude with discussions and conclusions in Sec. V.
@@ II.1 BLP measure
Breuer-Laine-Piilo Laine et al. (2010); Breuer et al. (2009) defined a measure that connects non-Markovianity with the information back-flow. This measure is important because it highlights the memory dependence on dynamics, which was not possible through divisibility-based non-Markovianity. It is not necessary to know the dynamical map to experimentally witness the non-Markovian behaviour of the dynamics. For defining the measure, they considered the trace distance defined as
@@
here, $P_{\rm dis}$ is the probability to distinguish two states. Unitary dynamics keeps this distance unchanged for all initial pairs; hence, distinguishability is unchanged under unitary dynamics. This is consistent with the fact that the system does not interact with any other quantum system under unitary dynamics, so there is no exchange of information. In the event of interactions based on distinguishability, there are two categories of dynamics: (a) Markovian dynamics: the distance does not increase for any initial pair for any interval of time. In this case, there is never an increase in distinguishability. (b) Non-Markovian dynamics: the distance increases at least for a few initial pairs for some time interval. In this case, the distinguishability may increase for an interval of time.
CP-divisible dynamics is CP for all intermediate initial to intermediate final times; this will keep contracting for all intermediate times since CP maps contract. P-divisible dynamics is CP from the initial to the final time, but for some intermediate times, it can be just P, not CP; it will also continue to contract because P-maps also contract. Because of this, CP-divisible and P-divisible dynamics are considered Markovian according to this measure Chruściński et al. (2018).
@@ II.2 Generalized measure
Thus, the BLP definition does not address the translation aspect of the dynamics. If two maps have the same $T$ part with different $\vec{c}$, then BLP can not differentiate them. To overcome this, the GBLP measure, discussed below, was introduced. The Helstorm Matrix
$\Delta_{p}(\rho_{1},\rho_{2})=p\rho_{1}-q\rho_{2}$ (4)
with $p+q=1$ was considered for the distance to overcome this issue, as it is sensitive to translations Chruściński et al. (2011). The generalized distance is defined as
@@
which is also a distance measure between two states, but prepared with biased probabilities. This is also shown by connecting the distinguishability of two states with the generalized distance Chruściński et al. (2011); Wißmann et al. (2015). Unlike the BLP measure, P-divisibility is known to be connected with this generalized measure with the if and only if condition.
As a simple example where BLP fails but GBLP works, consider the non-P-divisible generalized amplitude damping noise given by the following affine map $\mathcal{E}_{t}(\vec{r})=T(t)\vec{r}+\vec{c}(t)$, where
$T(t)$ $=\begin{pmatrix}\sqrt{\eta(t)}&0&0\\ 0&\sqrt{\eta(t)}&0\\ 0&0&\eta(t)\end{pmatrix}$
$\vec{c}(t)$ $=\begin{pmatrix}0\\ 0\\ (2s(t)-1)(1-\eta(t))\end{pmatrix}$ (6)
where, $\eta(t)=e^{-\gamma t}$ for $\gamma>0$ and $s(t)=\cos^{2}{ft}$ for all $f$. Consider the initial states $\vec{r}_{1}\equiv(x_{1},y_{1},z_{1})^{\dagger}$ and $\vec{r}_{2}\equiv(x_{2},y_{2},z_{2})^{\dagger}$. We have the trace distance between their evolved versions to be
$\frac{1}{2}\textrm{Tr}[\mathcal{E}_{t}(\rho_{1})-\mathcal{E}_{t}(\rho_{2})]$ $=\frac{1}{2}|T(t)\vec{r}_{1}-T(t)\vec{r}_{2}|$
@@
$\textrm{Tr}[p\mathcal{E}_{t}(\rho_{1})-q\mathcal{E}_{t}(\rho_{2})]$ $\neq|pT(t)\vec{r}_{1}-qT(t)\vec{r}_{2}+(p-q)\vec{c}(t)|$ (8)
@@
Figure 1: generalized distance between states $\dfrac{1}{2}(\mathbb{I}+\sigma_{z})$ and $\dfrac{1}{2}(\mathbb{I}+\sigma_{x})$ for $p=0.25$ is plotted as a function of time for generalized amplitude damping noise (See Eq. II.2). Noise parameters $\gamma$ and $f$ are 0.1 and 4, respectively. Distance and time are in arbitrary units.
@@ III Unitality, divisibility and distance
The above represents one example where the two measures differ; however, we aim to identify the general classes in which they differ. We are checking the distance evolution for both BLP and GBLP measures for CP, P and non-P maps. Once that is understood, the connection with the whole dynamics is straightforward because, for divisible dynamics, the whole dynamics is nothing but the composition of the intermediate maps. The GBLP measure is calculated for two initial states as $\rho_{i}\equiv(x_{i},y_{i},z_{i})$, where $i=1,2$ and probability $p$. The eigenvalues of $\Delta^{\dagger}\Delta$ are
@@
To calculate the distance, we add the square root of the eigenvalues, as the trace does not depend on the basis. The generalized distance is
@@
$D_{t}=\begin{array}[]{ll}\lvert\vec{w}(t)\rvert,&\lvert\vec{w}(t)\rvert\geq\lvert p-q\rvert\\ \lvert p-q\rvert,&\lvert\vec{w}(t)\rvert<\lvert p-q\rvert\\ \end{array}$ (9)
If $p=q$, the GBLP distance reverts to the BLP distance.
@@
The GBLP and BLP conditions are equivalent for qubit unital dynamics.
@@
The translation vector vanishes for unital dynamics, and the dynamics is characterised by rotation, compression or expansion. Eq. (9) reduces to:
$D_{t}=\begin{array}[]{ll}\lvert T(t)(p\vec{r}_{1}-q\vec{r}_{2})\rvert,&\lvert\vec{w}(t)\rvert\geq\lvert p-q\rvert\\ \lvert p-q\rvert,&\lvert\vec{w}(t)\rvert<\lvert p-q\rvert\\ \end{array}$ (10)
@@
    If the channel is GBLP NM, then $D_{t}(p\vec{r}_{1},q\vec{r}_{2})>D_{t-\delta}(p\vec{r}_{1},q\vec{r}_{2})$ for some $p$, $\vec{r}_{1}$, $\vec{r}_{2}$, $\delta$ and $t$. Since $p,q\leq 1$, $p\vec{r}_{1}$ and $q\vec{r}_{2}$ will also be valid vectors belonging to the Bloch sphere. Hence same distance can be written in terms of $\vec{r^{\prime}}_{1}$ and $\vec{r^{\prime}}_{2}$ where $\vec{r^{\prime}}_{1}=p\vec{r}_{1}$ and $\vec{r^{\prime}}_{2}=q\vec{r}_{2}$. The condition $D_{t}(p\vec{r}_{1},q\vec{r}_{2})>D_{t-\delta}(p\vec{r}_{1},q\vec{r}_{2})$ will become $D_{t}(\vec{r^{\prime}}_{1},\vec{r^{\prime}}_{2})>D_{t-\delta}(\vec{r^{\prime}}_{1},\vec{r^{\prime}}_{2})$ which is the condition for BLP-NM.
@@
    If the channel is BLP NM, then $D_{t}(\vec{r}_{1},\vec{r}_{2})>D_{t-\delta}(\vec{r}_{1},\vec{r}_{2})$ for some $\vec{r}_{1}$, $\vec{r}_{2}$, $\delta$ and $t$. This condition gives one pair that belongs to GBLP with $p=\dfrac{1}{2}=q$. That makes GBLP also NM.
As a corollary of Theorem 1, there is no unital non-P-divisible BLP Markovian dynamics. We now consider the question of what are the channels for which the BLP and GBLP criteria are nonequivalent.
@@
In the case of qubit dynamics, BLP and GBLP differ only for non-unital non-P-divisible BLP dynamics.
@@
Our earlier example demonstrated that the BLP and GBLP conditions are not equivalent. From Theorem 1, it follows that, nevertheless, they are equivalent for unital noise. Thus, their non-equivalence can hold only for non-unital noise. Here, by Ref. Chruściński et al. (2018), we know that P-divisibility is equivalent to the GBLP Markovianity. Thus, the non-equivalence of the BLP and GBLP conditions can arise only from non-unital non-P-divisible noise. If this noise is BLP non-Markovian, then the two criteria will agree. ∎